Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-18T18:59:53-0400

1)Let "(a,b)" be an open interval. Take arbitrary "x\in(a,b)". We have "B_\delta(x)\subset(a,b)",where "\delta=\min\{x-a,b-x\}", so "x" is an interior point of "(a,b)".

Since we take arbitrary "x\in(a,b)", we obtain that every point of "(a,b)" is an interior point of "(a,b)", so "(a,b)" is an open set.

2)Let "[a,b]" be a closed interval. Take arbitrary "x\in\mathbb R\setminus[a,b]". We have "B_\delta(x)\subset\mathbb R\setminus[a,b]", where "\delta=\min\{|x-a|,|b-x|\}", so "x" is an interior point of "\mathbb R\setminus[a,b]".

Since we take arbitrary "x\in\mathbb R\setminus[a,b]" , we obtain that every point of "\mathbb R\setminus[a,b]" is an interior point of "\mathbb R\setminus[a,b]" , so "\mathbb R\setminus[a,b]" is an open set. By the definition of closed set we have "[a,b]=\mathbb R\setminus(\mathbb R\setminus[a,b])" is a closed set.